# Just Shift It: Test-Time Prototype Shifting for Zero-Shot Generalization with Vision-Language Models

### v1

@@ Abstract
Advancements in vision-language models (VLMs) have propelled the field of computer vision, particularly in the zero-shot learning setting. Despite their promise, the effectiveness of these models often diminishes due to domain shifts in test environments. To address this, we introduce the Test-Time Prototype Shifting (TPS) framework, a pioneering approach designed to adapt VLMs to test datasets using unlabeled test inputs. Our method is based on the notion of modulating per-class prototypes in the shared embedding space. By pre-computing and caching prototypes generated with the pre-trained text encoder, TPS not only facilitates optimization-free prototype reuse for subsequent predictions but also enables seamless integration with current advancements in prompt engineering. At test-time, TPS dynamically learns shift vectors for each prototype based solely on the given test sample, effectively bridging the domain gap and enhancing classification accuracy. A notable aspect of our framework is its significantly reduced memory and computational demands when compared to conventional text-prompt tuning methods. Extensive evaluations across 15 datasets involving natural distribution shifts and cross-dataset generalization demonstrate TPS’s superior performance, achieving state-of-the-art results while reducing resource requirements. Code is available at https://github.com/elaine-sui/TPS.
Keywords:
test-time adaptation vision-language models zero-shot
@@ 1 Introduction
In recent years, the field of computer vision has witnessed remarkable progress, largely fueled by the emergence of robust vision-language foundation models [38, 9]. These models, upon undergoing pre-training on large-scale datasets, acquire a deep understanding of visual concepts, enabling their seamless application to a range of downstream tasks without task-specific training. While these foundation models exhibit much better zero-shot generalization abilities compared to ImageNet pre-trained models, they still suffer from performance degradation due to domain shifts at test-time.
To address this problem, prior studies have explored various fine-tuning techniques, ranging from traditional full-model tuning to more parameter-efficient methods such as tuning additional adapters [16, 8] or input prompts [22, 20]. Nevertheless, these strategies necessitate having enough labeled data for effective fine-tuning, posing challenges in domains where acquiring labeled data is difficult. In light of this, we turn our attention to the emerging paradigm of test-time adaptation (TTA), where model parameters are adjusted in an unsupervised manner with unlabeled test inputs.
@@
How can we effectively enable test-time adaptation using these advanced vision-language models (VLMs)? There have been relatively few initiatives in this area. A notable example is the Test-Time Prompt Tuning (TPT) method [43], which suggests fine-tuning several prompt tokens of the text input for each individual unlabeled test sample. However, this approach faces practical constraints due to the significant memory and computational demands involved in backpropagating through the text encoder for each image.
@@
In this work, we propose Test-Time Prototype Shifting (TPS), a simple yet effective framework that specifically adjusts per-class prototypes within the embedding space. Initially, we compute each class prototype using the pre-trained text encoder from a VLM, enabling the prototypes to be cached and reused for all subsequent predictions. At test-time, we adapt by learning a shift vector for each prototype on the fly for a single test sample, bridging the domain gap between the prototypes and the target sample. A key highlight of our framework is that the shift vectors are the only parameters being updated and are adjusted within the embedding space itself, circumventing the need for backpropagation through the text and visual encoders. In comparison to TPT, our TPS framework achieves a 10 times increase in speed while necessitating less than 1/10 of the memory cost.
@@
Benefiting from the two-stage learning paradigm, our approach fully capitalizes on the advancements in prompt engineering. Prototypes can be generated using vanilla templates such as “a photo of a {class}.", or crafted using more sophisticated, well-designed prompts [38, 52, 30] that significantly enhance generalization capabilities. Furthermore, experiments with few-shot learning on ImageNet reveal that learnable prompts are highly effective in addressing natural domain shifts [56, 57]. In our study, we build upon these previous advancements by generating improved prototypes that can be seamlessly integrated into our Test-Time Prototype Shifting framework.
@@
The zero-shot generalization capabilities of TPS were extensively evaluated across two distinct series of datasets: those involving natural distribution shifts and those focused on cross-dataset generalization. Our TPS consistently outperforms CLIP baselines, and surpasses current state-of-the-art (SoTA) by 3.3% on the natural distribution shifts benchmark and 1.9% on the cross-dataset generalization benchmark, respectively. Moreover, we demonstrate that regardless of the prototype-generation approach, learning a feature-space shift on the prototypes consistently boosts performance over zero-shot CLIP by over 4% on natural distribution shifts and up to 1% on cross-dataset generalization benchmarks. Remarkably, our approach not only out-performed TPT in terms of top-1 accuracy but also achieved this with only 1/10 of the memory and time costs.
@@
    Our TPS framework is designed to seamlessly integrate with existing advancements in prompt engineering, transforming it into a flexible plug-and-play system.
@@
    We achieve state-of-the-art performance on both natural distribution shift and cross-dataset generalization benchmarks, surpassing current SoTA by 3.3% and 1.9%, respectively. Additionally, our approach significantly reduces computational and memory demands by more than 10 times compared to TPT.
@@ 2.1 Test-Time Adaptation
Test-time adaptation (TTA) is the task of adapting a model’s weights on an unlabeled out-of-distribution test set in order to achieve higher test performance. In the context of vision tasks, traditional methods leverage ImageNet-pretrained image classifiers and use techniques such as computing pseudo-prototype class representations to update the linear classifier [18], learning better feature representations through self-supervised auxiliary tasks [46, 26, 24], adapting the normalization layers to learn the statistics of the target distribution [33, 32, 41, 44, 49], as well as minimizing prediction entropy to increase the confidence of predictions [49, 33, 53, 32, 10]. With the development of CLIP [38], recent work in TTA has been predominantly based on prompt tuning. This involves learning a tuneable text and/or image prompt to encode the visual distribution shift while maintaining the strong performances of these foundation models by keeping the pre-trained parameters fixed [43, 12, 28, 7, 55]. Despite only tuning a relatively small number of prompt parameters, tuning the input requires backpropagating through their respective encoders, which is especially memory intensive with large input sizes, making it infeasible in practice. Recently, an image diffusion-based method DiffusionTTA [37] has been proposed for TTA, involving updating classifier weights by training for the auxiliary task of image reconstruction using a conditional diffusion model. Nevertheless, this also necessitates backpropagation through the classifier and diffusion model. In contrast, this work proposes to avoid backpropagation through the encoders and maintain the richness of the CLIP embedding space by directly modulating the features within it.
Current TTA methods also comprise those that are considered training-free [11, 47]. Approaches include adding a parameter-free attention module to modulate multi-modal features [11] and computing the similarity between the target image and those from a constructed support set [47]. Although our work does not directly fit this setting, we follow the same spirit of minimally adjusting intermediate representations to close the domain gap.
@@ 3.2.1 Feature-Space Shift
The Test-Time Prompt Tuning (TPT) method, as proposed by Shu et al. [43], necessitates backpropagation through the text encoder at every test-time training step to learn the text prompt vectors. This process results in substantial computational and memory demands, which significantly hinders its practical applicability. Essentially, prompt tuning, when used for test-time adaptation, acts as an indirect technique for adjusting text embeddings. Its primary goal is to bridge the domain gap, simultaneously leveraging the multi-modal embedded knowledge of CLIP. Such considerations bring us to a pivotal question: why not directly modulate class prototypes within the embedding space? This has inspired the development of our Shift Learner module, designed specifically for learning modifications to the text embeddings. This approach not only capitalizes on the exceptional quality of the CLIP representation space but also adeptly avoids the requirement for the computationally demanding gradient backpropagation through the text encoder.
Rather than indirectly altering embeddings through prompt tuning, our approach is to directly learn to shift class prototypes within the embedding space. This strategy enables us to preserve the overarching architecture of the CLIP embedding space, while adjusting the class prototypes for better alignment with out-of-distribution visual embeddings. Our key proposition involves learning shifts for specific class prototypes, as opposed to a universal shift applicable to all classes. We recognize that domain gaps often consist of both dataset-level shifts (for example, transitioning from natural images to sketches) and class-level distribution shifts (such as differences between indoor and outdoor settings for dogs). A universally applied prototype shift, although effective for dataset-level adjustments, lacks the nuanced expressiveness required for class-level corrections. This is because it alters each prototype identically, maintaining the relative distances between class prototypes even after test-time training.
On the contrary, class-specific shifts have the capability to accommodate both levels of distribution shifts, enhancing their generalizability across various target datasets. Although adopting class-specific shifts results in the number of tunable parameters increasing linearly with the number of classes, this added complexity is minimal compared to the computational requirements of computing and storing gradients for the sizable text encoder.
@@ 3.2.2 Advanced Prototype Generation
In our research, we concentrate on unsupervised test-time adaptation. Consequently, we operate under the assumption that test images are unavailable before they are individually encountered at test-time, and specific knowledge about the type of domain shift is absent. For fair comparisons, we refrain from using target-domain-specific prompts. For instance, we avoid employing prompts such as “a photo of a sketch {class} exclusively for ImageNet Sketch, as this would imply prior knowledge of the domain shift type, which is contrary to our testing conditions. In this work, we employ three distinct prompt techniques: class-agnostic prompts using the set of 80 hand-crafted templates provided by CLIP [38] for ImageNet, class-specific prompts derived from intricate class descriptions generated by GPT-4 [34], and learnable prompts such as CoOp [56] and CoCoOp [57].
Aggregating class-specific representations. Our framework requires a single representation per class, but if such a representation is derived from a single prompt, the amount of information that it carries is limited to the number of input tokens that the CLIP text encoder was trained with. Hence, it would be best to have multiple class-specific representations, each providing complementary information in order to build a holistic representation of that class. Following [38, 30], we can easily improve the robustness of class prototypes by taking the mean of the class-specific embeddings. This allows us to leverage multiple prompting and prompt-learning techniques and retain the knowledge from these various representations. Further, this allows us to maintain the computational and memory efficiency of our method without sacrificing prototype representation strength. We combine the class-agnostic prompts and per-class descriptors to generate the final prototype set $\{\mathbf{p}_{c}\}_{c\in\mathcal{C}}$. Several types of combinations are discussed in the Appendix.
@@
1:pre-trained and frozen image encoder $f_{\mathcal{V}}$ from a VLM, pre-computed class prototypes $\mathbf{p}_{c}$ and trainable shift parameters $\mathbf{s}_{c}$ for $c\in\mathcal{C}$, test image $v_{0}$, set of augmentations $\mathcal{A}$, number of augmentations $(n-1)$
@@
11:     Update $\{\mathbf{s}_{c}\}_{c\in\mathcal{C}}$
@@ 3.2.3 Test-Time Training and Inference
At test time, given a single test image $v_{0}$, we follow [43] to augment it $(n-1)$ times and compute the features of the original and augmented images with the CLIP image encoder to obtain embeddings $\{\mathbf{x}_{i}\}_{i=0}^{n-1}$. As introduced in Section 3.2.1, we shift the pre-cached prototypes $\{\mathbf{p}_{c}\}_{c\in\mathcal{C}}$ to obtain $\{\mathbf{p}^{\prime}_{c}\}_{c\in\mathcal{C}}$. For each image feature $\mathbf{x}_{i}$, the predicted probabilities are calculated as
@@
where $\tau$ denotes the temperature scalar. Similar to TPT [43], we select the $k$ distributions with highest confidence (i.e. lowest entropy) of the batch and take the average. Denoting the image embeddings corresponding to the selected $k$ distributions as $\{\mathbf{x}^{\prime}_{i}\}_{i=1}^{k}$, we train our model to minimize the following entropy of this marginal distribution,
@@
where $\text{where }\tilde{p}(c|\mathbf{x}_{0})$ $=\frac{1}{k}\sum_{i=1}^{k}p(c|\mathbf{x}^{\prime}_{i},\mathbf{p}^{\prime}_{c})$ (4)
@@ 4.3.2 Natural Distribution Shifts
TPT [43]
68.98
54.77
63.45
77.06
47.94
62.44
60.81
TPT [43] + descriptors*
67.71
54.28
61.24
73.39
46.09
60.54
58.75
TPT + (templates + descriptors)*
69.54
55.13
63.95
77.46
48.34
62.88
61.22
DiffTPT [7]
70.30
55.68
65.10
75.00
46.80
62.28
60.52
Diffusion-TTA [37]
63.8
Ours (Shift + templates + descriptors) ${}_{(\uparrow{4.71})}$ 60.61${}_{(\uparrow{12.82})}$ 64.91${}_{(\uparrow{4.02})}$ 80.20${}_{(\uparrow{6.21})}$ 50.88${}_{(\uparrow{4.76})}$ 65.61${}_{(\uparrow{6.51})}$ 64.15${}_{(\uparrow{6.95})}$
@@
Zero-Shot Baseline
CLIP-ViT-B/16 + CoOp [56]
71.51
49.71
64.20
75.21
47.99
61.72
59.28
Test-Time Adaption Baselines
TPT + CoOp ([43, 56])
73.61
57.95
66.83
77.27
49.29
64.99
62.83
DiffTPT + CoOp ([7, 56]) 75.00
58.09
66.80
73.90
49.50
64.12
61.97
Ours (Shift + CoOp[56]) 73.73${}_{(\uparrow{2.22})}$ 60.49${}_{(\uparrow{10.78})}$ 66.84${}_{(\uparrow{2.64})}$ 77.44${}_{(\uparrow{2.23})}$ 49.08${}_{(\uparrow{1.09})}$ 65.52${}_{(\uparrow{3.80})}$ 63.46${}_{(\uparrow{4.18})}$

### v2

@@ Abstract
Advancements in vision-language models (VLMs) have propelled the field of computer vision, particularly in the zero-shot learning setting. Despite their promise, the effectiveness of these models often diminishes due to domain shifts in test environments. To address this, we introduce the Test-Time Prototype Shifting (TPS) framework, a pioneering approach designed to adapt VLMs to test datasets using unlabeled test inputs. Our method is based on the notion of modulating per-class prototypes in the shared embedding space. By pre-computing and caching prototypes generated with the pre-trained text encoder, TPS not only facilitates optimization-free prototype reuse for subsequent predictions but also enables seamless integration with current advancements in prompt engineering. At test-time, TPS dynamically learns shift vectors for each prototype based solely on the given test sample, effectively bridging the domain gap and enhancing classification accuracy. A notable aspect of our framework is its significantly reduced memory and computational demands when compared to conventional text-prompt tuning methods. Extensive evaluations across 15 image classification datasets involving natural distribution shifts and cross-dataset generalization, as well as in context-dependent visual reasoning, demonstrate TPS’s superior performance, achieving state-of-the-art results while reducing resource requirements. Code is available at https://github.com/elaine-sui/TPS.
@@ 1 Introduction
In recent years, the field of computer vision has witnessed remarkable progress, largely fueled by the emergence of robust vision-language foundation models [42, 10]. These models, upon undergoing pre-training on large-scale datasets, acquire a deep understanding of visual concepts, enabling their seamless application to a range of downstream tasks without task-specific ${}_{\!}$ While these foundation models exhibit much better zero-shot generalization abilities compared to ImageNet pre-trained models, they still suffer from performance degradation due to domain shifts at test-time.
To address this problem, prior studies have explored various fine-tuning techniques, ranging from traditional full-model tuning to parameter-efficient methods [17, 9, 26, 22]. Nevertheless, these strategies necessitate having enough labeled data for effective fine-tuning, posing challenges in domains where acquiring labeled data is difficult. In light of this, we turn our attention to the emerging paradigm of test-time adaptation (TTA), where model parameters are adjusted in an unsupervised manner with unlabeled test inputs.
@@
How can we effectively enable test-time adaptation using these advanced vision-language models (VLMs)? There have been relatively few initiatives in this area. A notable example is the Test-Time Prompt Tuning (TPT) method [47], which suggests fine-tuning several prompt tokens of the text input for each individual unlabeled test sample. However, this approach faces practical constraints due to the significant memory and computational demands involved in backpropagating through the text encoder for each image, as illustrated in Figure 1.
@@
In this work, we propose Test-Time Prototype Shifting (TPS), a simple yet effective framework that specifically adjusts per-class prototypes within the embedding space. Initially, we compute each class prototype using the pre-trained text encoder from a VLM, enabling the prototypes to be cached and reused for all subsequent predictions. At test-time, we adapt by learning a shift vector for each prototype on the fly for a single test sample, bridging the domain gap between the prototypes and the target sample. A key highlight of our framework is that the shift vectors are the only parameters being updated and are adjusted within the embedding space itself, circumventing the need for backpropagation through large encoders. In comparison to TPT, our TPS framework achieves a 10x increase in speed while necessitating less than 1/10 of the memory cost.
@@
Benefiting from the two-stage learning paradigm, our approach fully capitalizes on the advancements in prompt engineering, such as using more sophisticated, well-designed prompts that significantly enhance generalization capabilities [42, 55, 34], and using few-shot learned prompts [60, 61]. In our study, we build upon these previous advancements by generating improved prototypes that can be seamlessly integrated into our Test-Time Prototype Shifting framework.
@@
The zero-shot generalization capabilities of TPS were extensively evaluated across image classification and context-dependent visual reasoning. Image classification was evaluated with two distinct series of datasets: those involving natural distribution shifts and those focused on cross-dataset generalization. Our TPS consistently outperforms CLIP baselines, and surpasses current state-of-the-art (SoTA) by 3.3% on the natural distribution shifts benchmark and 1.9% on the cross-dataset generalization benchmark, respectively. Moreover, we demonstrate that regardless of the prototype-generation approach, learning a feature-space shift on the prototypes consistently boosts performance over zero-shot CLIP by over 4% on natural distribution shifts and up to 1% on cross-dataset generalization benchmarks. We further adapt TPS to context-dependent visual reasoning, outperforming current SoTA TTA methods on the Bongard-HOI benchmark [23] by 1.3%, on average. Remarkably, our approach not only outperformed TPT in terms of top-1 accuracy but also achieved this with only 1/10-th of the memory and time costs.
@@
    Our TPS framework seamlessly integrates with existing advancements in prompt engineering, transforming it into a flexible plug-and-play system.
@@
    We achieve state-of-the-art performance on image classification on both natural distribution shift and cross-dataset generalization benchmarks, surpassing current SoTA by 3.3% and 1.9%, respectively, as well as context-dependent visual reasoning, surpassing SoTA TTA methods by 1.3%. Additionally, our approach significantly reduces computational and memory demands by more than 10 times compared to TPT.
@@ 2.1 Test-Time Adaptation
Test-time adaptation (TTA) is the task of adapting a model’s weights on an unlabeled out-of-distribution test set in order to achieve higher test ${}_{\!}$ In the context of vision tasks, traditional methods leverage ImageNet-pretrained image classifiers and use techniques such as computing pseudo-prototype class representations to update the linear classifier [20], learning better feature representations through self-supervised auxiliary tasks [50, 30, 28], adapting the normalization layers to learn the statistics of the target distribution [37, 36, 45, 48, 52], as well as minimizing prediction entropy to increase the confidence of predictions [52, 37, 56, 36, 11]. With the development of CLIP [42], TTA methods have been predominantly based on prompt tuning. This involves learning a tunable text and/or image prompt to encode the visual distribution shift while maintaining the strong performances of these foundation models by keeping the pre-trained parameters fixed [47, 13, 32, 8, 59]. Despite only tuning a relatively small number of prompt parameters, tuning the input requires backpropagating through their respective encoders, which is especially memory intensive with large input sizes, making it infeasible in practice. Other works have proposed updating weights based on an auxiliary task [41], but nevertheless, requires backpropagation through the model. In contrast, this work proposes to avoid backpropagation through the encoders and maintain the richness of the CLIP embedding space by directly modulating the features in it.
Recent work have proposed training-free methods that directly modify the final predicted logit distribution. Some methods operate in an online streaming setting, using memory banks to store information on prior test inputs  [24, 58, 18]. This differs from our work where we assume that no prior knowledge of the test set is kept when predicting the label of each test example. Other training-free methods include adding a parameter-free attention module to modulate multi-modal features [12] and computing the similarity between the target image and those from a constructed support set [51]. Although our work does not directly fit this setting, we follow the same spirit of minimally adjusting intermediate representations to close the domain gap.
@@ 3.2.1 Feature-Space Shift
The Test-Time Prompt Tuning (TPT) method, as proposed by Shu et al. [47], necessitates backpropagation through the text encoder at every test-time training step to learn the text prompt vectors. This process results in substantial computational and memory demands, which significantly hinders its practical applicability. Essentially, prompt tuning, when used for test-time adaptation, acts as an indirect technique for adjusting text embeddings. Its primary goal is to bridge the domain gap, simultaneously leveraging the multi-modal embedded knowledge of CLIP. Such considerations bring us to a pivotal question: why not directly modulate class prototypes within the embedding space?
This inspired the development of our Shift Learner module, designed specifically for learning modifications to the text embeddings. This approach not only capitalizes on the exceptional quality of the CLIP representation space but also adeptly avoids the requirement for the computationally demanding gradient backpropagation through the text encoder. Rather than indirectly altering embeddings through prompt tuning, our approach is to directly learn to shift class prototypes within the embedding space. This strategy enables us to preserve the overarching architecture of the CLIP embedding space, while adjusting the class prototypes for better alignment with out-of-distribution visual embeddings.
@@ 3.2.2 Advanced Prototype Generation
${}_{\!}$ ${}_{\!}$ ${}_{\!}$ Our framework requires a single representation per class, but if such a representation is derived from a single prompt, the amount of information that it carries is limited to the number of input tokens that the CLIP text encoder was trained with. Hence, following [42, 34], we easily improve the robustness of class prototypes by taking the mean of the class-specific embeddings. This allows us to leverage multiple prompting and prompt-learning techniques and retain the knowledge from these various representations while keeping the computational and memory efficiency of our method without sacrificing prototype representation strength. We combine the class-agnostic prompts and per-class descriptors to generate the final prototype set $\{\mathbf{p}_{c}\}_{c\in\mathcal{C}}$. Several types of combinations are discussed in the Appendix.
@@
1:pre-trained and frozen image encoder $f_{\mathcal{V}}$ from a VLM, pre-computed class prototypes $\mathbf{p}_{c}$ and trainable shift parameters $\mathbf{s}_{c}$ for $c\in\mathcal{C}$, test image $v_{0}$, set of augmentations $\mathcal{A}$, number of augmentations $(n-1)$, optimizer Opt
@@
11:     Update $\{\mathbf{s}_{c}\}_{c\in\mathcal{C}}$: $\mathbf{s}_{c}\leftarrow\mathbf{s}_{c}-\texttt{Opt}(\partial\mathcal{L})$
@@ 3.2.3 Test-Time Training and Inference
At test time, given a single test image $v_{0}$, we follow [47] to augment it $(n-1)$ times and compute the features of the original and augmented images with the CLIP image encoder to obtain embeddings $\{\mathbf{x}_{i}\}_{i=0}^{n-1}$, where $\mathbf{x}_{0}$ is the embedding of the original image $v_{0}$, and $\{\mathbf{x}_{i}\}_{i=1}^{n-1}$ are the embeddings of the $(n-1)$ augmented versions of $v_{0}$. As introduced in Section 3.2.1, we shift the pre-cached prototypes $\{\mathbf{p}_{c}\}_{c\in\mathcal{C}}$ to obtain $\{\mathbf{p}^{\prime}_{c}\}_{c\in\mathcal{C}}$. For each image feature $\mathbf{x}_{i}$, the predicted probabilities are calculated as
@@
where $\tau$ denotes the temperature scalar. Similar to TPT [47], we select the $k$ distributions with highest confidence (i.e. lowest entropy) of the batch to filter out uninformative and misleading views of the image, and compute the average distribution from the selected $k$ distributions. Denoting the image embeddings corresponding to the selected $k$ distributions as $\{\mathbf{x}^{\prime}_{i}\}_{i=1}^{k}$, we train our model to minimize the following entropy of this marginal distribution,
@@
where $\text{where }\tilde{p}(c|\mathbf{x}_{0},\mathbf{p}^{\prime}_{c})$ $=\frac{1}{k}\sum_{i=1}^{k}p(c|\mathbf{x}^{\prime}_{i},\mathbf{p}^{\prime}_{c})$ (4)
This objective is used to encourage the model to make consistent, high-confidence predictions across multiple views, as model accuracy tends to correlate with model confidence [56, 47].
@@
3.2.4 TPS for Other Visual Tasks
Though we demonstrate how our test-time prototype shifting (TPS) method can work for image classification, it can also be used in more complex tasks such as Context-Dependent Visual Reasoning. In this work, we adapt the method by Shu et al. [47] by simply substituting the prompt-tuning step with our proposed shift-tuning step. We refer the reader to this paper for more details of the method.
@@ 4.1.1 Natural Distribution Shifts
Zero-Shot Baseline
CLIP-ViT-B/16 66.74 47.79 60.89 73.99 46.12 59.10 57.20
Test-Time Adaption Baselines
TPT [47] 68.98 54.77 63.45 77.06 47.94 62.44 60.81
TPT [47] + descriptors* 67.71 54.28 61.24 73.39 46.09 60.54 58.75
TPT + (templates + descriptors)* 69.54 55.13 63.95 77.46 48.34 62.88 61.22
DiffTPT [8] 70.30 55.68 65.10 75.00 46.80 62.28 60.52
Diffusion-TTA [41] 63.8 - - - - - -
Ours (Shift + templates + descriptors) 71.45(↑4.71) 60.61(↑12.82) 64.91(↑4.02) 80.20(↑6.21) 50.88(↑4.76) 65.61(↑6.51) 64.15(↑6.95)
@@
TPT + CoOp ([47, 60]) 73.61 57.95 66.83 77.27 49.29 64.99 62.83
DiffTPT + CoOp ([8, 60]) 75.00 58.09 66.80 73.90 49.50 64.12 61.97
Ours (Shift + CoOp[60]) 73.73(↑2.22) 60.49(↑10.78) 66.84(↑2.64) 77.44(↑2.23) 49.08(↑1.09) 65.52(↑3.80) 63.46(↑4.18)